Arianna invested $1,100 in an account paying an interest rate of 5.3% compounded daily. Assuming no deposits or withdrawals are made, how much money, to the nearest cent, would be in the account after 14 years?
Answer by Theo(13342)

assuming 365 days in a year, the future value will be $2,310.020303.

i used the Texas Instruments Business Analyst II calculator.

inputs were:

present value = -1100
future value = 0
payment = 0
number of time periods = 14 * 365 = 5110
interest rate per time period = 5.3 / 365.014520548%

i had the calculator compute the future value and the calculator told me the future value was $2,310.020303.

you can do the same with the online financial calculator at https://arachnoid.com/finance/

here's the display of the results.

$$$